# Mirror Descent Policy Optimisation for Robust Constrained Markov Decision Processes

### v1

@@ Conclusion
This paper presents mirror descent policy optimisation for robust constrained MDPs (RCMDPs), making use of policy gradient techniques to optimise both the policy and the transition kernel (as an adversary) on the Lagrangian representing a constrained MDP. In the oracle-based RCMDP setting, we confirm it is indeed possible to obtain guarantees similar to those of traditional MDPs, with an $\mathcal{O}\left(\frac{1}{T}\right)$ convergence rate for the squared distance as a Bregman divergence, and an $\mathcal{O}\left(e^{-T}\right)$ convergence rate for entropy-regularised objectives. In the sample-based setting, we require $\tilde{\mathcal{O}}(\epsilon^{-3})$ samples for an average regret of at most $\epsilon$, confirming an $\tilde{\mathcal{O}}\left(\frac{1}{T^{1/3}}\right)$ convergence rate. Experiments confirm the performance benefits of mirror descent policy optimisation in practice, obtaining significant improvements in test performance.
@@ Appendix A Bregman divergence and associated policy definitions
$\ell_{1}$-norm $\left\lVert p\right\rVert_{1}$ 00 for $x,y\in\mathbb{R}^{d}_{+}$.
@@
Parametrisation Update rule Uncertainty sets Entropy PTK: $p(s^{\prime}|s,a)=\frac{\bar{p}(s^{\prime}|s,a)\exp\left(\frac{\zeta^{\intercal}\phi(s^{\prime})}{\mathbf{\lambda^{\prime}}^{\intercal}\varphi(s,a)}\right)}{\sum_{s^{\prime\prime}}\bar{p}(s^{\prime\prime}|s,a)\exp\left(\frac{\zeta^{\intercal}\phi(s^{\prime\prime})}{\mathbf{\lambda^{\prime}}^{\intercal}\varphi(s,a)}\right)}$ $\xi_{t+1}\leftarrow\mathrm{arg}\max_{\xi}\{\left\langle\eta_{t}\nabla_{\xi}V_{\pi,p}(d_{0}),\xi\right\rangle-D(\xi,\xi_{t})\}$ $\mathcal{U}_{\xi}=\{\xi:D(\xi,\bar{\xi})\leq\kappa_{\xi}\}$. Gaussian mixture PTK: $p(s^{\prime}|s,a)=\sum_{m=1}^{M}\omega_{m}\mathcal{N}(\eta^{\intercal}\zeta(s,a))$ $\xi_{t+1}\leftarrow\mathrm{arg}\max_{\xi}\{\left\langle\eta_{t}\nabla_{\xi}V_{\pi,p}(d_{0}),\xi\right\rangle-D(\xi,\xi_{t})\}$ $\mathcal{U}_{\eta}=\{\eta:\forall m\in[M]\,D(\eta,\bar{\eta_{m}})\leq\kappa_{\eta}\}$ $\mathcal{U}_{\omega}=\{\omega\in\Delta^{M}:\forall m\in[M]\,\omega_{m}\in[F^{-1}_{\omega_{m}}(\delta/2),F^{-1}_{\omega_{m}}(1-\delta/2)]\}$
Appendix B Supporting lemmas for traditional MDPs
The proof relies on the following continuity and smoothness conditions.
Definition 4 (Continuity and smoothness).
A function $f:\Theta\to\mathbb{R}$ is $L_{\theta}$-Lipschitz continuous if $\left\lVert f(\theta)-f(\theta^{\prime})\right\rVert\leq L_{\theta}\left\lVert\theta-\theta^{\prime}\right\rVert$. $f$ is $l_{\theta}$-smooth if its gradients are $l_{\theta}$-Lipschitz continuous, i.e. $\left\lVert\nabla f(\theta)-\nabla f(\theta^{\prime})\right\rVert\leq l_{\theta}\left\lVert\theta-\theta^{\prime}\right\rVert$ or, equivalently,
@@ Appendix B Supporting lemmas for traditional MDPs
$|f(\theta^{\prime})-f(\theta)-\left\langle\nabla f(\theta),\theta^{\prime}-\theta\right\rangle|\leq\frac{l_{\theta}}{2}\left\lVert\theta-\theta^{\prime}\right\rVert_{2}^{2}\,.$
B.1 Softmax policy gradient with unregularised objective
Mei et al., (2020) previously used the above smoothness and continuity to prove the convergence rate for softmax policies on the unregularised objective. In particular, Theorem 4 in Mei et al., (2020) provides convergence rate results for softmax policies. We rephrase the theorem in a slightly more general manner based on the relation between the learning rate $\eta$ and the smoothness coefficient $l_{\theta}$, allowing easy reuse for our setting.
Theorem 5 (Convergence rate for softmax policies, Theorem 4 in Mei et al., (2020) rephrased).
Let $p$ be the transition dynamics, let costs range in $[0,1]$, let $\{\theta_{t}\}_{t\geq 1}$ be a sequence generated by the gradient descent update over logits according to
$\theta_{t+1}=\theta_{t}-\eta\nabla V_{\pi_{\theta_{t}},p}(\rho)\,,$ (58)
with $\eta=1/l_{\theta}$, and let $U_{p}=\inf_{s\in\mathcal{S},t\geq 1}\pi_{\theta}(a^{*}(s|p)|s)>0$. Then, for all $t\geq 1$, the regret is bounded by

### v2

@@ Conclusion
This paper presents mirror descent policy optimisation for robust constrained MDPs (RCMDPs), making use of policy gradient techniques to optimise both the policy and the transition kernel (as an adversary) on the Lagrangian representing a constrained MDP. In the sample-based RCMDP setting, we require $\tilde{\mathcal{O}}(\epsilon^{-3})$ samples for an average regret of at most $\epsilon$, confirming an $\tilde{\mathcal{O}}\left(\frac{1}{T^{1/3}}\right)$ convergence rate. An approximate algorithm for transition mirror ascent is also proposed which is of independent interest for designing adversarial environments in general MDPs. Experiments confirm the performance benefits of mirror descent policy optimisation in practice, obtaining significant improvements in test performance.
@@ Appendix A Bregman divergence and associated policy definitions
$\ell_{1}$-norm $\left\lVert p\right\rVert_{1}$ 0 for $x,y\in\mathbb{R}^{d}_{+}$.
@@
Parametrisation Update rule Uncertainty sets Entropy PTK: $p(s^{\prime}|s,a)=\frac{\bar{p}(s^{\prime}|s,a)\exp\left(\frac{\zeta^{\intercal}\phi(s^{\prime})}{\mathbf{\lambda^{\prime}}^{\intercal}\varphi(s,a)}\right)}{\sum_{s^{\prime\prime}}\bar{p}(s^{\prime\prime}|s,a)\exp\left(\frac{\zeta^{\intercal}\phi(s^{\prime\prime})}{\mathbf{\lambda^{\prime}}^{\intercal}\varphi(s,a)}\right)}$ $\xi_{t+1}\leftarrow\mathrm{arg}\max_{\xi}\{\left\langle\eta_{t}\nabla_{\xi}V_{\pi,p}(d_{0}),\xi\right\rangle-D(\xi,\xi_{t})\}$ $\mathcal{U}_{\xi}=\{\xi:D(\xi,\bar{\xi})\leq\kappa_{\xi}\}$. Gaussian mixture PTK: $p(s^{\prime}|s,a)=\sum_{m=1}^{M}\omega_{m}\mathcal{N}(\eta^{\intercal}\zeta(s,a))$ $\xi_{t+1}\leftarrow\mathrm{arg}\max_{\xi}\{\left\langle\eta_{t}\nabla_{\xi}V_{\pi,p}(d_{0}),\xi\right\rangle-D(\xi,\xi_{t})\}$ $\mathcal{U}_{\eta}=\{\eta:\forall m\in[M]\,,D(\eta,\bar{\eta}_{m})\leq\kappa_{\eta}\}$ $\mathcal{U}_{\omega}=\{\omega\in\Delta^{M}:\forall m\in[M]\,,\omega_{m}\in[F^{-1}_{\omega_{m}}(\delta/2),F^{-1}_{\omega_{m}}(1-\delta/2)]\}$
Appendix B Robust Sample-based PMD-PD pseudocode
B.1 Discretised state-action space
Algorithm 2 Robust Sample-based PMD-PD (discrete setting)
Inputs: Discount $\gamma\in[0,1)$, error tolerances $\epsilon,\epsilon_{0}^{\prime}>0$, failure probability $\delta\in[0,1)$, learning rates $\eta=\frac{1-\gamma}{\alpha}$, $\eta_{p}>0$, $\eta_{\lambda}=1.0$, and penalty coefficients $\alpha=\frac{2\gamma^{2}m\eta_{\lambda}}{(1-\gamma)^{3}}$ and $\alpha_{p}>0$.
Initialise: $\pi_{0}$ as uniform random policy, $\lambda_{0,j}=\max\left\{0,-\eta_{\lambda}\hat{V}_{\pi_{0},p_{0}}(\rho)\right\}$, $p_{0}=\bar{p}$.
for $k\in\left\{0,\dots,K-1\right\}$ do
  for $t\in\left\{0,\dots,t_{k}-1\right\}$ do
   for $(s,a)\in\mathcal{S}\times\mathcal{A}$ do
     Estimate the value: e.g. for softmax parametrisation,
   end for
   $\triangleright$ Policy mirror descent
   if Using Softmax parametrisation then
     $\pi_{k}^{t+1}(a|s)\leftarrow(\pi_{k}^{t}(a|s))^{1-\frac{\eta\alpha}{1-\gamma}}\exp\left(-\eta\frac{\hat{\mathbf{Q}}_{\pi_{k}^{t},p_{k}}^{\alpha}(s,a)}{1-\gamma}\right)\,~\forall(s,a)\in\mathcal{S}\times\mathcal{A}$.
   else
     $\triangleright$ Use general Bregman divergence
@@ B.1 Discretised state-action space
     $\theta_{k}^{t+1}\leftarrow\mathrm{arg}\min_{\theta\in\Theta}\eta\left\langle\nabla_{\theta}\hat{\mathbf{Q}}_{\pi_{k}^{t},p_{k}}^{\alpha},\theta\right\rangle+\alpha B_{d_{\rho}^{\pi_{\theta},p_{k}}}(\theta,\theta_{k})$
     Define $\pi_{k}^{t+1}:=\pi_{\theta_{k}^{t+1}}$.
   end if
  end for
  Define $\pi_{k+1}:=\pi_{k}^{t_{k}}$
  Apply Tabular Approximate TMA (Algorithm 1) with $c=\mathbf{c}_{k}$, $\pi=\pi_{k+1}$, and $p^{0}=p_{k}^{0}$ to get $p_{k}^{t_{k}^{\prime}}$.
  Define $p_{k+1}:=p_{k}^{t_{k}^{\prime}}$
  $\triangleright$ Augmented update of Lagrangian multipliers
  Generate $M_{V,k+1}$ samples of length $N_{V,k+1}$ based on $\pi_{k+1}$ and $p_{k+1}$ starting from $\rho$ ( Lemma 9).
end for
B.2 Continuous state-action space
Algorithm 3 Robust Sample-based PMD-PD (continuous setting)
Inputs: Discount factor $\gamma\in[0,1)$, sample sizes $M_{Q,k},M_{V,k}$ and episode lengths $N_{Q,k},N_{V,k}$ for all $k\in[K]$, learning rate $\eta$, TMA learning rate $\eta_{p}>0$, dual learning rate $\eta_{\lambda}$, error tolerance for LTMA $\epsilon_{0}^{\prime}>0$, and penalty coefficients $\alpha>0$ and $\alpha_{p}>0$.
Initialise: $\pi_{0}$ as uniform random policy, $\lambda_{0,j}=\max\left\{0,-\eta_{\lambda}\hat{V}_{\pi_{0},p_{0}}(\rho)\right\}$, $p_{0}=\bar{p}$.
for $k\in\left\{0,\dots,K-1\right\}$ do
  for $t\in\left\{0,\dots,t_{k}-1\right\}$ do
   Generate $M_{Q,k}$ samples of length $N_{Q,k}$ based on $\pi_{k}^{t}$ and $p_{k}$.
   Update $\hat{\mathbf{Q}}_{\pi_{k}^{t},p_{k}}\leftarrow\mathrm{arg}\min_{Q\in\mathcal{F}_{Q}}\frac{1}{M_{Q,k}}\sum_{j=1}^{M_{Q,k}}\left(Q(s_{0}^{j},a_{0}^{j})-\sum_{l=0}^{N_{Q,k}-1}\gamma^{l}\left[\tilde{\mathbf{c}}_{k}(s_{l}^{j},a_{l}^{j})\right]\right)^{2}$.
   $\triangleright$ Policy mirror descent
   $\theta_{k}^{t+1}\leftarrow\mathrm{arg}\min_{\theta\in\Theta}\eta\left\langle\nabla_{\theta}\hat{\mathbf{Q}}_{\pi_{k}^{t},p_{k}}^{\alpha},\theta\right\rangle+\alpha B_{d_{\rho}^{\pi_{\theta},p_{k}}}(\theta,\theta_{k})$.
   Define $\pi_{k}^{t+1}:=\pi_{\theta_{k}^{t+1}}$
  end for
  Define $\pi_{k+1}:=\pi_{k}^{t_{k}}$
  Apply Function Approximation TMA (Algorithm 1) with $c=\mathbf{c}_{k}$, $\pi=\pi_{k+1}$, and $p^{0}=p_{k}^{0}$ to get $p_{k}^{t_{k}^{\prime}}$.
  Define $p_{k+1}:=p_{k}^{t_{k}^{\prime}}$.
  $\triangleright$ Dual update of augmented Lagrangian multipliers
  Generate $M_{V,k+1}$ samples of length $N_{V,k+1}$ based on $\pi_{k+1}$ and $p_{k+1}$ starting from $\rho$.
end for
Appendix C Supporting lemmas and proofs for Approximate TMA
C.1 Recursion property
The following straightforward recursion property is useful in the proof of the regret bound of approximate TMA.
Lemma 11 (Recursion property, Lemma 13 in Xiao, (2022)).
Let $\beta\in(0,1)$, and $b>0$. If $\{a_{t}\}_{t\geq 0}$ is a non-negative sequence such that $x_{t+1}\leq\beta a_{t}+b$, then $x_{t}\leq\beta^{t}x_{0}+\frac{b}{1-\gamma}$.
C.2 Regret bound for approximate TMA
The regret bound of Approximate TMA (Theorem 1) is shown below.
Proof.
Note that
$\langle p^{*}(\cdot|s,a)-p^{t+1}(\cdot|s,a),\hat{G}_{\pi,p^{t}}(s,a,\cdot)\rangle=\langle p^{*}(\cdot|s,a)-p^{t}(\cdot|s,a),\hat{G}_{\pi,p^{t}}(s,a,\cdot)\rangle+\langle p^{t}(\cdot|s,a)-p^{t+1}(\cdot|s,a),\hat{G}_{\pi,p^{t}}(s,a,\cdot)\rangle\,.$
From the above and the pushback property in Eq. 27 with $p=p^{*}$,
$\langle p^{*}(\cdot|s,a)-p^{t}(\cdot|s,a),G_{\pi,p^{t}}(s,a,\cdot)\rangle+\langle p^{t}(\cdot|s,a)-p^{t+1}(\cdot|s,a),G_{\pi,p^{t}}(s,a,\cdot)\rangle\leq\frac{\alpha_{p}(t)}{\eta_{p}(t)}\left(B(p^{*},p^{t})-B(p^{*},p^{t+1})\right)$
Weighting both sides according to the occupancy measure, we obtain
$\underbrace{\frac{1}{1-\gamma}\sum_{s\in\mathcal{S}}d_{\rho}^{\pi,p^{*}}(s)\sum_{a\in\mathcal{A}}\pi(a|s)\langle p^{*}(\cdot|s,a)-p^{t}(\cdot|s,a),\hat{G}_{\pi,p^{t}}(s,a,\cdot)\rangle}_{A}+$
$\underbrace{\frac{1}{1-\gamma}\sum_{s\in\mathcal{S}}d_{\rho}^{\pi,p^{*}}(s)\sum_{a\in\mathcal{A}}\pi(a|s)\langle p^{t}(\cdot|s,a)-p^{t+1}(\cdot|s,a),\hat{G}_{\pi,p^{t}}(s,a,\cdot)\rangle}_{B}$
$\leq\frac{\alpha_{p}(t)}{\eta_{p}(t)}\left(B_{d_{\rho}^{\pi,p^{*}}}(p^{*},p^{t})-B_{d_{\rho}^{\pi,p^{*}}}(p^{*},p^{t+1})\right)\,.$
For part A), note that
$\frac{1}{1-\gamma}\sum_{s\in\mathcal{S}}d_{\rho}^{\pi,p^{*}}(s)\sum_{a\in\mathcal{A}}\pi(a|s)\langle p^{*}(\cdot|s,a)-p^{t}(\cdot|s,a),\hat{G}_{\pi,p^{t}}(s,a,\cdot)\rangle$
$\frac{1}{1-\gamma}\sum_{s\in\mathcal{S}}d_{\rho}^{\pi,p^{*}}(s)\sum_{a\in\mathcal{A}}\pi(a|s)\left(\langle p^{*}(\cdot|s,a)-p^{t}(\cdot|s,a),G_{\pi,p^{t}}(s,a,\cdot)\rangle+\langle p^{*}(\cdot|s,a)-p^{t}(\cdot|s,a),\hat{G}_{\pi,p^{t}}(s,a,\cdot)-G_{\pi,p^{t}}(s,a,\cdot)\rangle\right)$ (decomposing)
$\geq V_{\pi,p^{*}}(\rho)-V_{\pi,p^{t}}(\rho)+\frac{1}{1-\gamma}\sum_{s\in\mathcal{S}}d_{\rho}^{\pi,p^{*}}(s)\sum_{a\in\mathcal{A}}\pi(a|s)\langle p^{*}(\cdot|s,a)-p^{t}(\cdot|s,a),\hat{G}_{\pi,p^{t}}(s,a,\cdot)-G_{\pi,p^{t}}(s,a,\cdot)\rangle$ (Lemma 22)
$\geq V_{\pi,p^{*}}(\rho)-V_{\pi,p^{t}}(\rho)-\frac{2C\epsilon^{\prime}}{1-\gamma}\,$,
where the last step bounds the absolute value as in the proof of Lemma 5.
For part B), note that
$\frac{1}{1-\gamma}\sum_{s\in\mathcal{S}}d_{\rho}^{\pi,p^{*}}(s)\sum_{a\in\mathcal{A}}\pi(a|s)\langle p^{t}(\cdot|s,a)-p^{t+1}(\cdot|s,a),\hat{G}_{\pi,p^{t}}(s,a,\cdot)\rangle$
$=\frac{1}{1-\gamma}\sum_{s\in\mathcal{S}}d_{\rho}^{\pi,p^{*}}(s)\sum_{a\in\mathcal{A}}\pi(a|s)\big(\langle p^{t}(\cdot|s,a)-p^{t+1}(\cdot|s,a),G_{\pi,p^{t}}(s,a,\cdot)\rangle+$
$\hskip 170.71652pt\langle p^{t}(\cdot|s,a)-p^{t+1}(\cdot|s,a),\hat{G}_{\pi,p^{t}}(s,a,\cdot)-G_{\pi,p^{t}}(s,a,\cdot)\rangle\big)$ (decomposing)
$\frac{1}{1-\gamma}\sum_{s\in\mathcal{S}}\frac{d_{\rho}^{\pi,p^{*}}(s)}{d_{\rho}^{\pi,p^{t+1}}(s)}d_{\rho}^{\pi,p^{t+1}}(s)\sum_{a\in\mathcal{A}}\pi(a|s)\big(\langle p^{t}(\cdot|s,a)-p^{t+1}(\cdot|s,a),G_{\pi,p^{t}}(s,a,\cdot)\rangle+$
$\hskip 170.71652pt\langle p^{t}(\cdot|s,a)-p^{t+1}(\cdot|s,a),\hat{G}_{\pi,p^{t}}(s,a,\cdot)-G_{\pi,p^{t}}(s,a,\cdot)\rangle\big)$
$\geq\left\lVert\frac{d_{\rho}^{\pi,p^{*}}}{d_{\rho}^{\pi,p^{t+1}}}\right\rVert_{\infty}\Big(\left(V_{\pi,p^{t}}(\rho)-V_{\pi,p^{t+1}}(\rho)\right)+$
$\hskip 113.81102pt\frac{1}{1-\gamma}\sum_{s\in\mathcal{S}}d_{\rho}^{\pi,p^{t+1}}(s)\sum_{a\in\mathcal{A}}\pi(a|s)\langle p^{t}(\cdot|s,a)-p^{t+1}(\cdot|s,a),\hat{G}_{\pi,p^{t}}(s,a,\cdot)-G_{\pi,p^{t}}(s,a,\cdot)\rangle\Big)$ (ascent property (Eq. 28))
$\geq M\Big(\left(V_{\pi,p^{t}}(\rho)-V_{\pi,p^{t+1}}(\rho)\right)+$